# Foveated Retinotopy Improves Classification and Localization in Convolutional Neural Networks

### v4

@@ Acknowledgments
Authors received funding from the ANR project number ANR-20-CE23-0021 (“AgileNeuRobot”) and from the french government under the France 2030 investment plan, as part of the Initiative d’Excellence d’Aix-Marseille Université – A*MIDEX grant number AMX-21-RID-025 “Polychronies”. This work was granted access to the HPC resources of Aix-Marseille Université financed by the project Equip@Meso (ANR-10-EQPX-29-01) of the program “Investissements d’Avenir” supervised by the Agence Nationale de la Recherche. For the purpose of open access, authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v5

@@ Acknowledgments
Authors received funding from the ANR project number ANR-20-CE23-0021 (AgileNeuRobot) and from the french government under the France 2030 investment plan, as part of the Initiative d’Excellence d’Aix-Marseille Université – A*MIDEX grant number AMX-21-RID-025 Polychronies.
This work was granted access to the HPC resources of Aix-Marseille Université financed by the project Equip@Meso (ANR-10-EQPX-29-01) of the program “Investissements d’Avenir” supervised by the Agence Nationale de la Recherche.
For the purpose of open access, the authors have applied a CC‑BY public copyright licence to any Author‑Accepted Manuscript version arising from this submission.
Author Contributions
All authors contributed equally to the following roles according to the CRediT Taxonomy: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Resources, Data Curation, Writing - Original Draft, Writing - Review & Editing, and Visualization. LP and ED provided Supervision and LP performed Project Administration, and Funding Acquisition. All authors have read and agreed to the published version of the manuscript.
Competing Interests
We affirm that this manuscript is original, has not been published, and is not under consideration elsewhere. All authors have read and approved the final version of the manuscript. We have adhered to ethical research guidelines and declare no conflicts of interest.